Cheyenne Construction Company has entered into a contract beginning January 1, 2017, to build a parking complex. It has been estimated that the complex will cost $596,000 and will take 3 years to construct. The complex will be billed to the purchasing company at $895,000. The following data pertain to the construction period.

2017

2018

2019

Costs to date $268,200 $482,760 $605,000
Estimated costs to complete 327,800 113,240 –0–
Progress billings to date 267,000 552,000 895,000
Cash collected to date 237,000 502,000 895,000


(a) Using the percentage-of-completion method, compute the estimated gross profit that would be recognized during each year of the construction period. (If answer is 0, please enter 0. Do not leave any fields blank.)

Gross profit recognized in 2017 $

Gross profit recognized in 2018 $

Gross profit recognized in 2019 $


(b) Using the completed-contract method, compute the estimated gross profit that would be recognized during each year of the construction period. (If answer is 0, please enter 0. Do not leave any fields blank.)

Gross profit recognized in 2017 $

Gross profit recognized in 2018 $

Gross profit recognized in 2019 $
Answer #1
Revenue recognition as per Percentage of Completion Method - Cheyenne Cosntruction Company
Year Actual cost incurred (A) Total cost incurred till date (B) Total estimated Cost ( C ) % of completion (D= B/C) Contract Price
(E )
Total Revenue to be recoganized
(F=E * D)
Revenue for current period (G) Gross Profit (H=G-A)
2017 $268,200 $268,200 $596,000 45.00% $895,000 $402,750 $402,750 $134,550
(268200+327800)
2018 $214,560 $482,760 $596,000 81.00% $895,000 $724,950 $322,200 $107,640
(482760+113240) (724950-402750)
2019 $122,240 $605,000 $605,000 100.00% $895,000 $895,000 170050 $47,810
(895000-724950)
Revenue recognition as per Completed Contract method
Year Actual cost incurred (A) Total cost incurred till date (B) Total estimated Cost ( C ) Contract Price Gross Profit
(D)
2017 $268,200 $268,200 $596,000 $895,000 $0
2018 $214,560 $482,760 $596,000 $895,000 $0
2019 $122,240 $605,000 $605,000 $895,000 $290,000.00